# A Statistical Framework for Understanding Causal Effects that Vary by Treatment Initiation Time in EHR-based Studies

### v1

@@ Abstract
Comparative effectiveness studies using electronic health records (EHR) consider data from patients who could “enter” the study cohort at any point during an interval that spans many years in calendar time. Unlike treatments in tightly controlled trials, real-world treatments can evolve over calendar time, especially if comparators include standard of care, or procedures where techniques may improve. Efforts to assess whether treatment efficacy itself is changing are complicated by changing patient populations, with potential covariate shift in key effect modifiers. In this work, we propose a statistical framework to estimate calendar-time specific average treatment effects and describe both how and why effects vary across treatment initiation time in EHR-based studies. Our approach projects doubly robust, time-specific treatment effect estimates onto candidate marginal structural models and uses a model selection procedure to best describe how effects vary by treatment initiation time. We further introduce a novel summary metric, based on standardization analysis, to quantify the role of covariate shift in explaining observed effect changes and disentangle changes in treatment effects from changes in the patient population receiving treatment. Extensive simulations using EHR data from Kaiser Permanente are used to validate the utility of the framework, which we apply to study changes in relative weight loss following two bariatric surgical interventions versus no surgery among patients with severe obesity between 2005–2011.
@@ 1 Introduction
Electronic health record (EHR) databases provide a useful data source for comparative effectiveness studies, allowing for larger sample sizes from patient populations which are more diverse than those typical of clinical trials (Casey2016EHR). In EHR-based studies, study entry periods frequently extend over a decade or more (arterburn2020, fisher2018association, coleman2016longterm, Harrington2024, haneuse2025, o2013long), perhaps because (a) treatment initiation is uncommon, (b) long-term outcomes are of interest, and/or (c) extending data collection further back in time requires little additional effort. For example, in a large EHR-based study, arterburn2020 examined long-term weight loss outcomes of patients undergoing bariatric surgery between 2005-2015 in comparison to matched non-surgical controls. In contrast to treatment protocols in tightly controlled randomized clinical trials, real-world treatments, such as those examined by arterburn2020, can evolve over time, especially if comparators include a somewhat loosely defined standard of care (Laiteerapong2015) (e.g., for obesity management), or bariatric surgical procedures where techniques or best practices may change (Sheetz2020, Faria2017, Felsenreich2021, Guldogan2020Sleeve). As such, causal effects of interest in EHR-based studies with long study entry periods may vary over the course of treatment initiation time, which we also refer to as calender time throughout.
@@
Heterogeneity in treatment initiation time has been discussed previously in the context of studies in pharmacoepidemiology. Common across such studies, however, is that variation in treatment initiation time is most often viewed as a potential source of bias towards estimating an implicitly assumed constant treatment effect across calendar time (though for longitudinal or time-to-event outcomes, the effect may vary across time since initiation). For example, stratification of propensity score models by calendar year has been suggested as a solution to address changes in physician prescription patterns or characteristics of treated patients over time (Dusetzina2013, Mack2013, Dilokthornsakul2014CalendarTime). When comparisons of interest involve, say, two drugs approved at different points in time, analysis is often restricted to a period where both medications are available to reduce differences attributable to calendar time (Gagne2013TreatmentDynamics, Suissa2017PrevalentNewUser). Strategies based on instrumental variables (Gokhale2018CalendarTime) and matching (HelinSalmivaara2014SequentialCohort) have also been proposed. In all, the predominant viewpoint has been that variation in causal contrasts by calendar time is something that must be adjusted for, rather than adoption of causal estimands which explicitly use calendar time in their definition. That this is the case likely owes to the fact that in pharmacoepidemiologic studies, typical treatments (e.g., drugs or medications) are not changing over time. In such cases, there may be strong reason to believe there exists a common effect over calendar time in a static population, though the underlying populations may be shifting.
@@
Causal estimands which depend on calendar time have recently been explored in the context of stepped-wedge cluster randomized trials (wang2024model, Lee2025SteppedWedge) and staggered-adoption difference-in-difference settings (callaway2021did, sun2021dynamic), the latter of which often considers analysis of multi-year periods during which a policy is rolled out. Because intervention timing is planned or quasi-experimental across clusters, both study designs provide a natural way to define calendar time-specific effects for groups of clusters initiating treatment at the same time. By contrast, EHR-based studies analyze patient-level outcomes under naturally occurring treatment patterns, which do not follow any controlled schedule, and thus strategies from these literatures do not directly apply.
@@
Determining whether causal effects vary over calender time in EHR-based studies can help inform whether some fundamental aspect of treatment efficacy is changing, and whether researchers might report common or calendar time-varying effect estimates. However, efforts to assess whether treatment efficacy itself is changing are complicated by changing patient populations, and thus, potentially, covariate shift in key effect modifiers. As such, it is equally important to understand why causal effect estimates may vary over time, and the degree to which effect modification plays a role.
@@
In this work, we propose a statistical framework to estimate and describe how and why causal effects vary over calendar time in EHR-based studies. First, doubly-robust, calendar time-specific average treatment effects (ATE) are estimated on the basis of pooled nonparametric nuisance models. Second, estimates are projected onto candidate marginal structural models (MSM), each representing a different pattern in how effect estimates vary across treatment initiation time (e.g., constant, linear etc.). A model selection procedure is then applied to choose which MSM best describes temporal variation in effect estimates. Next, we define a causal estimand that corresponds to the treatment effect that would arise if the distribution of covariates among eligible patients treated at a given point in time instead reflected the distribution observed at some alternative point in time. Finally, we introduce a novel summary metric to quantify the role of covariate shift in explaining observed effect changes and disentangle changes in treatment effects from changes in the patient population receiving treatment. To our knowledge, this work is the first to apply tools from semiparametric theory (tsiatis_semiparametric_2006, kennedy2023semiparametric) to study calendar time-varying treatment effects in EHR-based studies.
@@
This work is motivated by the use of EHR data to study weight loss outcomes following bariatric surgery, particularly in comparison to similarly obese patients who do not undergo surgery, a setting which is described in detail in Section 2, and woven throughout the work. Section 3 formalizes notation and describes calendar time-specific causal estimands. Sections 4 and 5 introduce statistical tools, based on semiparametric efficiency theory, for determining how and why causal effects vary over calendar time, respectively, while Section 6 outlines how analysts can synthesize information from estimation and inferential procedures outlined in previous sections to determine what to report in their own study. Section 7 returns to the motivating study to apply the methods outlined in this work, and finally, Section 8 concludes with discussion. Detailed proofs and simulations which validate tools from this framework are provided in the Supplementary Material.
2 An EHR-based Study of Bariatric Surgery
@@
Bariatric surgery is an intervention to mitigate obesity and related comorbidities, with typical candidates having a body mass index (BMI) exceeding 35 kg/m2 (kaiserpermanente2023). Due to the cost and complexity of conducting large-scale trials for bariatric surgery (Courcoulas2024), EHR-based studies are a critical tool to study long-term outcomes following bariatric surgery, particularly in relation to similarly obese patients not undergoing surgery (arterburn2020, obrien2018microvascular, fisher2018association, Harrington2024, coleman2022bariatric). Despite its success as a weight-loss intervention, only 1-2% of eligible candidates elect to undergo bariatric surgery annually (arterburn2014bariatric). Therefore, EHR-based studies of bariatric surgery must consider a multi-year study entry period to obtain a meaningful number of patients undergoing surgery.
@@
The two most common bariatric surgical procedures are Roux-en-Y gastric bypass (RYGB) and sleeve gastrectomy (SG), with SG having surpassed RYGB in frequency in the past 15 years (arterburn2014bariatric, Reames2014). SG is a less invasive and technically easier procedure than RYGB, and as such is offered by more surgeons. Relevant to this work is that the surgical technique for SG has historically been somewhat less standardized than for RYGB, particularly during the time-frame during which SG became the more commonly performed procedure (Guldogan2020Sleeve, Rosenthal2012, Gagner2013). Given the rapid shifts in the relative popularity of RYGB and SG in recent years, along with possibly evolving surgical techniques and standards of care for obesity management, understanding potential temporal changes in the effectiveness bariatric surgery is of particular interest.
@@
With this backdrop, we examined relative weight change at four pre-specified times post-surgery (6 months, 1 year, 2 years, 3 years), comparing 17,905 patients undergoing bariatric surgery between 2005-2011 at one of three Kaiser Permanente (KP) sites (KP Washington, KP Northern California, KP Southern California) with 933,044 non-surgical patients. In particular, we compared outcomes between surgery and no surgery, each of RYGB/SG versus no surgery separately, and RYGB versus SG. We leveraged EHR data from DURABLE, an NIH-funded study examining long-term outcomes of bariatric surgery across Kaiser Permanente sites (arterburn2020, coleman2016longterm, obrien2018microvascular, coleman2022bariatric, fisher2018association, Harrington2024, haneuse2025).
@@
Using this data source, arterburn2020 previously characterized longitudinal weight trajectories following RYGB and SG in comparison to matched non-surgical patients, though calendar time was considered only as a potential confounder. As an alternative to the matched cohort design employed by arterburn2020, we used a sequential target trial emulation approach (benz2024, madenci2024, haneuse2025, cain2010, danaei2013observational, caniglia2023). Specifically, we considered a sequence of 84 target trials (formalized in Section 3), initiating at monthly intervals between January 2005 and December 2011. Following arterburn2020, our study population included adults aged 19-79 years with BMI 35 kg/m2. We excluded patients without a year of continuous enrollment in KP insurance and those pregnant in the year prior to the start of a particular target trial. 17,596 surgical patients (98.2%) and 763,016 non-surgical patients (81.8%) met the study inclusion criteria at baseline of at least one target trial.
@@
On the basis of these data, one might choose to estimate and report common a effect for bariatric surgery on 3 year relative weight change, in comparison to non-surgical patients. Though formal details are deferred until Section 7, such an approach produced estimates of -19.8% (-19.9%, -19.6%) for surgery overall, with corresponding estimates of -23.1% (-23.3%, -22.9%) for RYGB and -16.3% (-16.6%, -16.1%) and SG, in close agreement with estimates of -23.5% and -16.7% reported by arterburn2020, respectively. These summaries, however, do not acknowledge the potential for variation across calendar time, and thus do not offer any insight into possible temporal variation in the effectiveness of bariatric surgery.
Figure 1 summarizes the distribution of several key covariates among eligible bariatric patients across the 84 target trials (i.e., across calendar time). Notably, the distribution of several effect modifiers highlighted by arterburn2020 (age, BMI, race, and diabetes status) varies meaningfully across time of surgery. For example, the mean BMI of patients undergoing bariatric surgery in a given calendar month declines from 47.5 kg/m2 in January 2005 to 43.8 kg/m2 in December 2011, with monthly averages ranging between 42.9 kg/m2 and 49.6 kg/m2. Accordingly, any attempt to study how the effect of bariatric surgery may have changed over the course of the study period should acknowledge and account for this covariate shift. A similar figure displaying covariate shift for non-surgical patients is available in the Supplementary Materials (Figure S1).
@@ 3.1 Notation
When a treatment strategy does not correspond to a concrete clinical event, time zero can be challenging to define (hernan2016immortal). In order to ensure that the start of follow-up is well-defined for all patients, particularly those not undergoing surgery, we partition the study entry period into a sequence of $M$ target trials (benz2024, madenci2024, cain2010, danaei2013observational, caniglia2023), indexed by $m\in\{1,\dots,M\}$. In the motivating study of bariatric surgery, $M=84$ and component target trials are conducted at monthly intervals between January 2005 ($m=1$) and December 2011 ($m=M=84$).
To the best of our knowledge, the sequential target trial emulation design has been applied exclusively in applications examining time-to-event outcomes (scola2023implementation). Since relative weight change is not a time-to-event endpoint, we describe notation to formalize this design for continuous outcomes. Let $A_{m}\in\{0,1\}$ denote a binary treatment (e.g., surgery vs. no surgery) received at baseline of trial $m$, $Y_{m}\in\mathbb{R}$ denote an outcome of interest (e.g., relative weight change since baseline of trial $m$), and $\bm{L}_{m}\in\mathbb{R}^{p}$ denote a set of covariates at baseline of trial $m$ assumed sufficient for control of confounding, such as those in Figure 1. Let $E_{m}=g(\bm{L}_{m})$ be a binary eligibility indicator for inclusion in target trial $m$, where $g(\cdot)$ represents a fixed and known eligibility rule (benz2024, benz2025). Practically, because not all patients are in the EHR for the entire duration of the study entry period, we treat subjects as ineligible $(E_{m}=0)$ for periods before they appear and/or after they disappear from the EHR. Thus, at each of $M$ regularly occurring intervals, we ascertain $(\bm{L}_{m},A_{m},Y_{m})$ for eligible subject-trials; formally, we work with coarsened observed data units:
$O=(E_{1},E_{1}\bm{L}_{1},E_{1}A_{1},E_{1}Y_{1},\dots,E_{m},E_{m}\bm{L}_{m},E_{m}A_{m},E_{m}Y_{m},\dots,E_{M},E_{M}\bm{L}_{M},E_{M}A_{M},E_{M}Y_{M}).$
We observe a sample $(O_{1},\ldots,O_{n})$ of $n$ independent and identically distributed observations from an underlying and unknown data distribution, $P$, i.e., the joint distribution of information across both patients and trials. Implicitly throughout, we assume covariates across trials share a common domain $\mathcal{L}$. As will become apparent in Sections 3.2 and 4, covariate, treatment and outcome values of ineligible patients are not required for the definition of our calendar time-specific causal estimands, nor for estimation.
For any $P$-integrable function $f$ we define $Pf=\mathbb{E}_{P}[f]=\int f\,dP$ to denote the expectation of $f$ under $P$, and analogously define $\mathbb{P}_{n}f=\frac{1}{n}\sum_{i=1}^{n}f(O_{i})$ to express the sample mean of $f$ on data units $(O_{1},...,O_{n})$. Lastly, we let $Y_{m}(a_{m})$ denote counterfactual outcomes at time $m$, that is the outcome that would have been observed under treatment $A_{m}=a_{m}$ (hernan2024).

### v2

@@ Abstract
Standard practice in electronic health record (EHR)-based studies evaluating the comparative effectiveness of bariatric surgery relative to no surgery is to estimate and report a constant treatment effect across calendar time. However, real-world treatment strategies can evolve, particularly when comparators include standard of care or surgical procedures where techniques may improve, making it clinically important to ascertain whether efficacy of bariatric surgery has changed over time. Efforts to determine whether treatment efficacy itself is evolving are complicated by changing patient populations, with potential covariate shift in key effect modifiers. Through a comprehensive analysis of EHR data from Kaiser Permanente following two bariatric surgical procedures compared to standard of care, we develop a statistical framework to estimate calendar time-specific average treatment effects and describe both how and why effects vary across treatment initiation time in EHR-based studies. Our approach projects doubly robust, time-specific treatment effect estimates onto candidate marginal structural models and uses a model selection procedure to best describe how effects vary by treatment initiation time. We further introduce a novel summary metric, based on standardization analysis, to quantify the role of covariate shift in explaining observed effect changes and disentangle changes in treatment effects from changes in the patient population receiving treatment.
Keywords: causal inference, treatment effect heterogeneity, bariatric surgery, electronic health records
@@ 1 Introduction
Electronic health record (EHR) databases provide a useful data source for comparative effectiveness studies, allowing for larger sample sizes from patient populations which are more diverse than those typical of clinical trials. In EHR-based studies, study entry periods frequently extend over a decade or more (Arterburn et al., 2020, Fisher et al., 2018, Coleman et al., 2016, Harrington et al., 2024, Haneuse et al., 2025), perhaps because (a) treatment initiation is uncommon, (b) long-term outcomes are of interest, and/or (c) extending data collection further back in time requires little additional effort. For example, in a large EHR-based study, Arterburn et al. Arterburn et al. (2020) examined long-term weight loss outcomes of patients undergoing bariatric surgery between 2005-2015 in comparison to matched non-surgical controls. In contrast to treatment protocols in tightly controlled randomized clinical trials, real-world treatments, such as those examined by Arterburn et al. Arterburn et al. (2020), can evolve over time, especially if comparators include a somewhat loosely defined standard of care (Laiteerapong and Huang, 2015) (e.g., for obesity management), or bariatric surgical procedures where techniques or best practices may change (Sheetz et al., 2020, Faria, 2017, Felsenreich et al., 2021, Guldogan et al., 2020). As such, causal effects of interest in EHR-based studies with long study entry periods may vary over the course of treatment initiation time, which we also refer to as “calendar time” throughout.
@@
Heterogeneity in treatment initiation time has been discussed previously in the context of studies in pharmacoepidemiology. Common across such studies, however, is that variation in treatment initiation time is most often viewed as a potential source of bias towards estimating an implicitly assumed constant treatment effect across calendar time (though for longitudinal or time-to-event outcomes, the effect may vary across time since initiation). For example, stratification of propensity score models by calendar year has been suggested as a solution to address changes in physician prescription patterns or characteristics of treated patients over time (Dusetzina et al., 2013, Mack et al., 2013, Dilokthornsakul et al., 2014). When comparisons of interest involve, say, two drugs approved at different points in time, analysis is often restricted to a period where both medications are available to reduce differences attributable to calendar time (Gagne et al., 2013, Suissa et al., 2017). Strategies based on instrumental variables (Gokhale et al., 2018) and matching (Helin-Salmivaara et al., 2014) have also been proposed. In all, the predominant viewpoint has been that variation in causal contrasts by calendar time is something that must be adjusted for, rather than adoption of causal estimands which explicitly use calendar time in their definition. That this is the case likely owes to the fact that in pharmacoepidemiologic studies, typical treatments (e.g., drugs or medications) are not changing over time. In such cases, there may be strong reason to believe there exists a common effect over calendar time in a static population, though the underlying populations may be shifting.
@@
Causal estimands which depend on calendar time have recently been explored in the context of stepped-wedge cluster randomized trials (Wang et al., 2024, Lee et al., 2025) and staggered-adoption difference-in-difference settings (Callaway and Sant’Anna, 2021, Sun and Abraham, 2021), the latter of which often considers analysis of multi-year periods during which a policy is rolled out. Because intervention timing is planned or quasi-experimental across clusters, both study designs provide a natural way to define calendar time-specific effects for groups of clusters initiating treatment at the same time. By contrast, EHR-based studies analyze patient-level outcomes under naturally occurring treatment patterns, which do not follow any controlled schedule, and thus strategies from these literatures do not directly apply.
@@
Determining whether causal effects vary over calendar time in EHR-based studies is of clinical interest as such efforts can help characterize the real-world history of a treatment and inform whether some fundamental aspect of its efficacy is changing. Moreover, describing the structure of causal effects over calendar time yields an estimate of the treatment effect at the end of the study period (i.e., that which most closely reflects current practice) and informs how historical data might be leveraged to guide present-day clinical decision-making. However, such efforts are complicated by changing patient populations, and thus, potentially, covariate shift in key effect modifiers. As such, it is equally important to understand why causal effect estimates may vary over time, and the degree to which effect modification plays a role.
@@
In this work, motivated by our experience in a series of studies of long-term outcomes following bariatric surgery (Arterburn et al., 2020, Coleman et al., 2016, O’Brien et al., 2018, Coleman et al., 2022, Fisher et al., 2018, Harrington et al., 2024, Haneuse et al., 2025), we propose a statistical framework to estimate and describe how and why causal effects vary over calendar time in EHR-based studies. First, doubly-robust, calendar time-specific average treatment effects (ATE) are estimated on the basis of pooled nonparametric nuisance models. In an effort to summarize temporal trends, estimates are projected onto candidate marginal structural models (MSM), each representing a different pattern in how effect estimates vary across treatment initiation time (e.g., constant, linear, etc.). An MSM selection procedure is then applied to choose which structure best describes temporal variation in effect estimates. Next, to disentangle changes in treatment effects from changes in the patient population receiving treatment, we define a causal estimand that corresponds to the treatment effect which would have been observed if the covariate distribution among eligible patients at a given time instead reflected that at some alternative time. Finally, we introduce a metric summarizing variability across these “cross-temporal” contrasts to quantify the role of covariate shift in explaining observed effect changes across calendar time. To our knowledge, this work is the first to apply tools from semiparametric theory (Tsiatis, 2006, Kennedy, 2023) to study calendar time-varying treatment effects in EHR-based studies.
@@
As mentioned, this work is directly motivated by the use of EHR data to study weight loss outcomes following bariatric surgery, particularly in comparison to similarly obese patients who do not undergo surgery, a setting which is described in detail in Section 2, and woven throughout the work. Section 3 formalizes notation and describes calendar time-specific causal estimands. Sections 4 and 5 introduce statistical tools, based on semiparametric efficiency theory, for determining how and why causal effects vary over calendar time, respectively, while Section 6 outlines how analysts can synthesize information from estimation and inferential procedures outlined in previous sections to determine what to report in their own study. Section 7 presents results leveraging tools from this framework to detect and characterize calendar time-varying effects in our study of weight loss following bariatric surgery. Finally, Section 8 concludes with discussion. Detailed proofs and simulations which validate tools from this framework are provided in the Supplementary Material.
2 An EHR-based Study of Weight Loss Following Bariatric Surgery
2.1 Bariatric Surgery
@@
Bariatric surgery is an intervention to mitigate obesity and related comorbidities, with typical candidates having a body mass index (BMI) exceeding 35 kg/m2. Due to the cost and complexity of conducting large-scale trials for bariatric surgery (Courcoulas et al., 2024), EHR-based studies are a critical tool to study long-term outcomes following bariatric surgery, particularly in relation to similarly obese patients not undergoing surgery (Arterburn et al., 2020, O’Brien et al., 2018, Fisher et al., 2018, Harrington et al., 2024, Coleman et al., 2022). Despite its success as a weight-loss intervention, only 1-2% of eligible candidates elect to undergo bariatric surgery annually (Arterburn and Courcoulas, 2014). Therefore, EHR-based studies of bariatric surgery must consider a multi-year study entry period to obtain a meaningful number of patients undergoing surgery.
@@
The two most common bariatric surgical procedures are Roux-en-Y gastric bypass (RYGB) and sleeve gastrectomy (SG), with SG having surpassed RYGB in frequency in the past 15 years (Arterburn and Courcoulas, 2014). SG is a less invasive and technically easier procedure than RYGB, and as such is offered by more surgeons. Relevant to this work is that the surgical technique for SG has historically been somewhat less standardized than for RYGB, particularly during the time-frame during which SG became the more commonly performed procedure (Guldogan et al., 2020, Rosenthal et al., 2012, Gagner et al., 2013). Given the rapid shifts in the relative popularity of RYGB and SG in recent years, along with possibly evolving surgical techniques and standards of care for obesity management, understanding potential temporal changes in the effectiveness of bariatric surgery is of particular interest. Not only does characterization of this temporal structure offer insight into these procedures during an important change in their utilization history, but it also includes assessments of treatment effects at the end of the study period that are more relevant to current clinical practice when the data are proximal to present day.
2.2 The DURABLE EHR Database
@@
With this backdrop, we examined relative weight change at four pre-specified times post-surgery (6 months, 1 year, 2 years, 3 years), comparing 17,905 patients undergoing bariatric surgery between 2005-2011 at one of three Kaiser Permanente (KP) sites (KP Washington, KP Northern California, KP Southern California) with 933,044 non-surgical patients. In particular, we compared outcomes between surgery and no surgery, each of RYGB/SG versus no surgery separately, and RYGB versus SG. We leveraged EHR data from DURABLE, an NIH-funded study examining long-term outcomes of bariatric surgery across Kaiser Permanente sites (Arterburn et al., 2020, Coleman et al., 2016, O’Brien et al., 2018, Coleman et al., 2022, Fisher et al., 2018, Harrington et al., 2024, Haneuse et al., 2025).
@@
Using this data source, Arterburn et al. Arterburn et al. (2020) previously characterized longitudinal weight trajectories following RYGB and SG in comparison to matched non-surgical patients, though calendar time was considered only as a potential confounder. As an alternative to the matched cohort design employed by Arterburn et al. Arterburn et al. (2020), we used a sequential target trial emulation approach (Benz et al., 2025a, Madenci et al., 2024, Haneuse et al., 2025, Danaei et al., 2013, Caniglia et al., 2023). Specifically, we considered a sequence of 84 target trials (formalized in Section 3), initiating at monthly intervals between January 2005 and December 2011. Following Arterburn et al., our study population included adults aged 19-79 years with BMI exceeding 35 kg/m2. We excluded patients without a year of continuous enrollment in KP insurance and those pregnant in the year prior to the start of a particular target trial. In total, 17,596 surgical patients (98.2%) and 763,016 non-surgical patients (81.8%) met the study inclusion criteria at baseline of at least one target trial.
@@
2.3 Calendar Time-Varying Effects in Bariatric Surgery
Towards quantifying the relative benefits of surgery versus no surgery, the standard way forward would be to estimate and report a common effect across calendar time, for example, on 3-year relative weight change. Though formal details are deferred until Section 7, such an approach produces estimates of -19.8% (-19.9%, -19.6%) for surgery overall, with corresponding estimates of -23.1% (-23.3%, -22.9%) for RYGB and -16.3% (-16.6%, -16.1%) for SG. Note, these are in close agreement with estimates of -23.5% and -16.7% reported by Arterburn et al. Arterburn et al. (2020), respectively. While such summaries have appeal, particularly for communication to broad audiences, they may conceal meaningful changes in the effectiveness of bariatric surgery over calendar time, whether due to evolving surgical techniques, shifting patient populations, or perhaps both.
Towards gaining preliminary insight, Figure 1 summarizes the distribution of several key covariates among eligible bariatric patients across the 84 target trials (i.e., across calendar time). Notably, the distribution of several effect modifiers highlighted by Arterburn et al. Arterburn et al. (2020) (age, BMI, race, and diabetes status) varies meaningfully across calendar time. For example, mean BMI of patients undergoing bariatric surgery in a given calendar month declines from 47.5 kg/m2 in January 2005 to 43.8 kg/m2 in December 2011, with monthly averages ranging between 42.9 kg/m2 and 49.6 kg/m2. Accordingly, any attempt to study how the effect of bariatric surgery may have changed over the course of the study period should acknowledge and account for this covariate shift. A similar figure displaying covariate shift for non-surgical patients is available in the Supplementary Materials (Figure S1).
@@ 3 Problem Setting
In a hypothetical clinical trial, each patient would be randomized to undergo or not undergo surgery and be followed from the point of treatment assignment. By contrast, patients in our EHR-based study are not randomized to their respective treatment assignment. A key challenge in addressing the primary questions laid out in Section 2 is that, in contrast to patients who undergo surgery, there is no obvious way to assign a “time zero” to patients who are in the “no surgery” arm (Hernán et al., 2016). To resolve this dilemma, we partition the study entry period into a sequence of $M$ target trials (Benz et al., 2025a, Madenci et al., 2024, Danaei et al., 2013, Caniglia et al., 2023), indexed by $m\in\{1,\dots,M\}$. At baseline of each target trial, study eligibility is queried for all patients in the EHR, and patients are included in each trial for which they meet the eligibility criteria. In our study of bariatric surgery, $M=84$ and component target trials are conducted at monthly intervals between January 2005 ($m=1$) and December 2011 ($m=M=84$).
@@ 3.1 Notation
To the best of our knowledge, the sequential target trial emulation design has been applied exclusively in applications examining time-to-event outcomes (Scola et al., 2023). Since relative weight change is not a time-to-event endpoint, we describe notation to formalize this design for univariate continuous outcomes. Let $A_{m}\in\{0,1\}$ denote a binary point exposure (e.g., surgery vs. no surgery) received at baseline of trial $m$, $Y_{m}\in\mathbb{R}$ denote an outcome of interest (e.g., relative weight change since baseline of trial $m$), and $\bm{L}_{m}\in\mathbb{R}^{p}$ denote a set of covariates at baseline of trial $m$ assumed sufficient for control of confounding, such as those in Figure 1. Let $E_{m}=g(\bm{L}_{m})$ be a binary eligibility indicator for inclusion in target trial $m$, where $g(\cdot)$ represents a fixed and known eligibility rule (Benz et al., 2025a, b). Practically, because not all patients are in the EHR for the entire duration of the study entry period, we treat subjects as ineligible $(E_{m}=0)$ for periods before they appear and/or after they disappear from the EHR. Thus, at each of $M$ regularly occurring intervals, we ascertain for each patient $i$ their eligibility $E_{m,i}$ for trial $m$, and, if eligible, $\bm{Z}_{m,i}\coloneqq(\bm{L}_{m,i},A_{m,i},Y_{m,i})$. That is, patients contribute $\bm{Z}_{m,i}$ for all target trials where they meet study eligibility. Formally, we work with coarsened observed data units
$O_{i}=(E_{1,i},E_{1,i}\bm{Z}_{1,i},\dots E_{m,i},E_{m,i}\bm{Z}_{m,i},\dots E_{M,i},E_{M,i}\bm{Z}_{M,i})$
We observe a sample $(O_{1},\ldots,O_{n})$ of $n$ independent and identically distributed observations from an underlying and unknown data distribution, $P$, i.e., the joint distribution of information across both patients and trials. For succinctness, we often drop the patient index $i$ and work with a generic observation $O$. Implicitly throughout, we assume covariates across trials share a common support contained in $\mathcal{L}$. As will become clear in Sections 3.2 and 4, covariate, treatment and outcome values of ineligible patients are not required for the definition of our calendar time-specific causal estimands, nor for estimation.
For any $P$-integrable function $f$ we define $Pf=\mathbb{E}_{P}[f]=\int f\,dP$ to denote the expectation of $f$ under $P$, and analogously define $\mathbb{P}_{n}f=\frac{1}{n}\sum_{i=1}^{n}f(O_{i})$ to express the sample mean of $f$ on data units $(O_{1},...,O_{n})$. Lastly, we let $Y_{m}(a_{m})$ denote counterfactual outcomes for the $m^{\text{th}}$ trial, that is the outcome that would have been observed under treatment $A_{m}=a_{m}$ (Hernán and Robins, 2024).